# Holistic Tokenizer for Autoregressive Image Generation

### v1

@@ Abstract
The vanilla autoregressive image generation model generates visual tokens in a step-by-step fashion, which limits the ability to capture holistic relationships among token sequences. Moreover, most visual tokenizers map local image patches into latent tokens, leading to limited global information. To address this, we introduce Hita, a novel image tokenizer for autoregressive (AR) image generation. It introduces a holistic-to-local tokenization scheme with learnable holistic queries and local patch tokens. Besides, Hita incorporates two key strategies for improved alignment with the AR generation process: 1) it arranges a sequential structure with holistic tokens at the beginning followed by patch-level tokens while using causal attention to maintain awareness of previous tokens; and 2) before feeding the de-quantized tokens into the decoder, Hita adopts a lightweight fusion module to control information flow to prioritize holistic tokens. Extensive experiments show that Hita accelerates the training speed of AR generators and outperforms those trained with vanilla tokenizers, achieving 2.59 FID and 281.9 IS on the ImageNet benchmark. A detailed analysis of the holistic representation highlights its ability to capture global image properties such as textures, materials, and shapes. Additionally, Hita also demonstrates effectiveness in zero-shot style transfer and image in-painting. The code is available at https://github.com/CVMI-Lab/Hita.
@@ 1 Introduction
To address this, we introduce Hita, a novel global-to-local image tokenizer for AR image generation. First, we propose to utilize a set of learnable queries to capture the holistic information from the patch-level embeddings of an image. Meanwhile, a pre-trained foundation model, such as DINOv2 [7], is adopted to inject semantics into the learned queries to enhance their representation. Then, we utilize 2 separate codebooks to quantize the learned queries and patch-level embeddings, respectively, resulting in holistic tokens and patch-level tokens. Next, we fuse the de-quantized holistic and patch-level tokens together with a novel token fusion module to reconstruct the image. Such design can better prioritize the holistic token and improve both image reconstruction and generation quality.
Once Hita is trained, a vanilla autoregressive (AR) transformer model, e.g. Llama [48], can be seamlessly incorporated to generate both holistic and image patch tokens sequentially, which can be decoded into image facilitated by Hita. During the AR generation, the AR generator first generates holistic tokens, which encapsulate the global image information and serve as a prefix prompt. These holistic tokens then guide the subsequent generation of local image patch tokens, alleviating difficulties and maintaining global coherence to improve generation quality.
Our in-depth ablation experiments confirm that the learned holistic tokens effectively capture global image features, such as shape, color, texture, etc, from pixels (see Fig. 2a). This has been demonstrated through effortlessly enabling style transfer by combining holistic tokens from reference images with patch-level tokens from the source image. Moreover, without task-specific training, the model can be seamlessly applied to image inpainting by encoding the given incomplete image into latent tokens, which serve as prefix sequence prompts to the AR model, requiring it to generate visual tokens for completing the image (see Fig. 2b), demonstrating the capability of holistic tokens in ensuring global consistent synthesis.
@@
To summarize, our contribution can be listed as follow:
  • We propose a novel global-to-local image tokenizer for AR image generation, with which an AR generation model can be integrated without any modification.
  • The proposed image tokenizer emerges with new characteristics, such as zero-shot style stransfer and zero-shot image in-painting.
  • Extensive experiments conducted on class-conditional image generation task demonstrates it can accelerate the converge speed during training and significantly improve the synthesis quality.
@@ 2 Related Work
TiTok [59] and VAR [47] are related efforts but differ from Hita in several aspects. TiTok aims to create compact 1D tokens with queries to reduce redundancy, while Hita uses queries to capture holistic properties for 1D holistic tokens and fuses them with 2D image patch tokens to reconstruct images. TiTok refines queries for 1D tokens tailored for reconstruction objectives, whereas Hita extracts information from pre-trained foundation model and patch features to capture global information. TiTok requires a complex training strategy, while Hita enables efficient single-stage training. Hita outperforms TiTok in AR generation and offers new features like zero-shot style transfer and inpainting. VAR [47] predicts global and local tokens indiscriminately, while Hita separates holistic and patch-level tokens, enriching holistic tokens with pre-trained foundation to form global representations and uses causal attention to align with the causal nature of AR models, which VAR lacks. Removing initial coarse-scale tokens seldom affects VAR’s reconstruction, indicating they are not crucial.
@@ 3 Method
In this section, we first introduce the preliminary related to image tokenizer and AR generation models. Then, we present the design of our holistic tokenizer step by step. Next, their characteristics are also discussed.
@@ 3.1 Background
Autoregressive Image Generation. When used for visual generation, the 2D quantized map from a tokenizer is converted into a 1D sequence in a pre-defined order, e.g. raster-scan order. Then, an AR generation model predicts the sequence of image tokens autoregressively. To achieve this, AR models typically incorporate causal attention with prefix token sequence to predict the next, which focuses on learning the dependencies between the prefix discrete image tokens and the current one. Considering a sequence of discrete tokens ${x_{1:N}}$, an AR model is trained on these tokens by maximizing the cross-entropy depicted in Eq. (3), where ${c}$ is class label or text embedding, ${x_{j<i}}$ serves as the prefix sequence of tokens for the current token ${x_{i}}$.
@@
During inference, the AR model initially predicts visual tokens by sampling tokens from the learned distribution autoregressively. Subsequently, the generated visual tokens are converted back into images through the decoder $\mathcal{D}(\cdot)$.
@@ 3.2 Holistic Tokenization for AR Generation
As shown in Fig. 3, Hita incorporates the following key components. First, to capture holistic property, we introduce learnable queries that interact with image patch embeddings via attention modules. A pre-trained foundation model (e.g., DINOv2 [7]) injects semantic-level features. To align Hita’s latent space with the causal nature of AR models, we use a causal transformer to integrate holistic and patch features before quantization, embedding the sequential structure into the tokenization process to learn prefix-aware latent tokens. Details are depicted in Sec. 3.2.1. Then, the holistic and patch-level features are quantized using separate codebooks (see Sec. 3.2.2). Finally, in the token fusion and decoding procedure after de-quantization, Hita first integrates a fusion module with casual transformers to process these tokens in order and prioritize holistic tokens. Then they are fed into the decoder for image reconstruction (see Sec. 3.2.3).
@@ 3.2.1 Holistic Feature Extraction
We build Hita based on VQGAN [12]. Specifically, we adopt its encoder $\mathcal{E}(\cdot)$ to encode input image $I$ into local image patch embeddings, and further incorporate a set of learnable queries ${Q}$ to gather holistic information from the whole input. To enhance the holistic representation, a pre-trained foundation model $\mathcal{H}(\cdot)$ like DINOv2 [7] is leveraged to inject rich semantic-level features. To achieve this, we first concatenate the learnable queries and the flattened patch-level embeddings as well as foundation model features into sequence and feed them into a simple transformer $\mathcal{E}_{\text{{trans}}}(\cdot)$ is utilized, where holistic queries ${Q}$ interact with all patch-level image embeddings and the injected semantic-level features. Considering the AR model models the probability of observing the current token based on its predecessors, a well-aligned latent space with the causal nature is encouraged. Therefore, the extracted holistic feature and patch-level feature are concatenated and reorganized in an ordered sequence that holistic queries come first followed by the patch-level image patch feature obtained in raster-scan order. This sequence is then fed into a transformer with global features serving as prefix prior. The whole process is formulated as Eq. (4).
@@
Figure 4: The illustration of token fusion and decoding procedure
@@ 3.2.2 Separate Vector Quantization
Once both learned queries $\overline{Q}$ and patch-level images $\overline{Z}$ are produced, they are quantized with holistic quantizer $\mathcal{Q}_{H}(\cdot)$ and patch quantizer $\mathcal{Q}_{P}(\cdot)$, respectively. Then, the holistic tokens $\hat{Q}$ and patch-level $\hat{Z}$ tokens could be obtained. Considering the codebook significantly impacts the performance of the image tokenizer. Following [54, 46], we apply ${\ell}_{2}$-normalization to the codebook vectors, opting for a lower vector dimension and a larger codebook size to improve the reconstruction quality and codebook utilization.
@@ 3.2.3 Token Fusion and Decoding Procedure
After quantization, the holistic tokens and patch-level tokens could be fed into a vanilla transformer to perform token fusion and then fed into the decoder to reconstruct the image. However, such a design incurs codebook collapse in the holistic quantizer (3th row in Table. 3). This occurs because patch-level tokens directly impact the reconstruction of relevant image patches through the skip connection in the transformer, which makes it bypass holistic tokens (see Fig. 1a in Appendix). As a result, the token fusion module learned a trivial solution, which overlooks the holistic tokens and leads to holistic codebook collapse.
@@ 3.2.4 Training objective
3.3 Autoregresive Image Generation
Once the tokenizer is trained, a standard autoregressive (AR) transformer model, e.g. Llama [48], can be seamlessly integrated without requiring any revision to generate these latent tokens sequentially. Starting from the class/text embedding $c$, the AR model first generates holistic tokens as prefix prompts, then guides the subsequent generation of patch token sequence in the way of next-token prediction. Once all the visual tokens are generated, they will be tackled by the token fusion and decoding procedure to be converted into pixels. Additionally, in AR models, each visual token is added with positional embedding. Since holistic tokens form a 1D sequence, we add learnable positional embeddings for them. Meanwhile, image patch tokens continue to use 2D RoPE [45] for their positional embeddings, which remains the same as LlamaGen [46]
@@ 3.4 Discussions
What do holistic tokens capture? To understand what has been learned by the holistic tokens, we perform an experiment on the learned holistic tokens. Specifically, two images are first encoded and quantized into holistic tokens and image patch tokens, respectively. Then, the holistic tokens from one image are concatenated with the patch tokens from the other. These tokens are fed through the de-quantization process and the decoder to synthesize a new image. This entire procedure is conducted using a pre-trained holistic tokenizer, without any additional model training or fine-tuning. As shown in Fig. 2a, the holistic tokens can capture global information from an image, such as texture, color, materials, shapes, and other features from the pixels. For instance, in the case of the lizard image, Hita successfully transfers the material properties of the ship from the reference image in thumbnail to the input image, creating a realistic mechanical lizard. More samples are shown in Appendix.
Does holistic tokenizer enhance semantic coherence? To verify whether the holistic features can enhance the global coherence of the generated content, we perform a zero-shot image in-painting evaluation. In this in-painting task, the model is forced to complete the lower half of an image given only the upper half. This setup prevents local information leakage during in-painting. We first tokenize the images into holistic and patch tokens, then feed them to the AR model to generate additional tokens, conditioned on class embeddings in AR model. Hita-XXL is used without modifications to the model architecture or parameter tuning.
As shown in Fig. 2b, Hita effectively completes the lower part of the image based on the upper part, producing images that maintain strong semantic consistency. In contrast, outputs from LlamaGen [46] lack this consistency. For example, in the case of the fish image, Hita correctly generates a complete fish, while LlamaGen [46] mistakenly generates a “fish-bird”. These results confirm that our holistic tokenizer is semantic-aware and has a stronger ability to generate semantically and holistically coherent examples. Furthermore, the ablation study of linear probing on holistic tokens is conducted in Sec. 4.
@@ 4.1 Setup
Hita is optimized on the ImageNet [8] training set and evaluated on the validation set. To ensure a fair comparison with LlamaGen [46], we train the tokenizer on $336\times 336$ images, which yields a comparable number of discrete tokens to it. We also adopt the training settings for tokenizer training from LlamaGen. Additionally, during evaluation, the test images are resized to $256\times 256$, consistent with the evaluation procedure in LlamaGen.
@@
Following the AR generation procedure in LlamaGen [46], the AR models first generate images of the same size as the image tokenizer and then resize them to $256\times 256$ for evaluation. Unless otherwise specified, all models are trained with the same settings as LlamaGen, and the entire training process lasts for 50 epochs. Additionally, classifier-free guidance [18] is used in the AR generation process.
@@ 4.3 Ablation Study
In the following, we first study the impact of learnable queries and semantic injection. Secondly, we perform analysis on the selection length of $k$. Next, the different foundation model for semantics injection is discussed. Later, the semantic level of holistic tokens is also studied. Further ablations are also provided in the Appendix.
Learnable queries and Semantic Injection Analysis. In Hita, it incorporates a set of learnable queries and leverages semantic injection from the pre-trained model (e.g., DINOv2 [7]) to capture holistic information. To validate their effectiveness, we conduct component analysis on the learnable queries ${Q}$ and semantic injection $\mathcal{H}$. Specifically, We first train a tokenizer with only the proposed attention modules as the baseline. Then, we gradually add learnable queries and a pre-trained foundation model (e.g., DINOv2) to the baseline. Next, we train a Hita-B AR generation model for 50 epochs to estimate the generation quality. Moreover, we conduct linear probing on the whole latent tokens after the last attention module in the baseline. As for the comparison experiments, we only perform linear probing on the holistic tokens from the last attention module. As depicted in Table. 2, simply introducing learnable queries can significantly improve image reconstruction and generation quality. The additional features from the foundation model can further enhance the reconstruction and generation quality. Besides, as indicated by the linear probing, the learnable queries can capture better semantic information compared to the baseline. With semantics injected, this semantic-level representation can be further enhanced. This illustrates that the combination of the learnable queries and semantic injection is essential for learning better holistic information, as lacking either degrades image reconstruction and generation quality.
@@
Is $k\textgreater 0$ crucial? To emphasize holistic tokens and avoid its codebook collapse, Hita first builds a 1D ordered sequence with holistic tokens first and followed by the flattened patch-level tokens. Next, it is fed into a causal transformer for token fusion. Subsequently, the last $k$ holistic tokens will combined with patch-level tokens to reconstruct the image. To verify the doubt, we first configure $k$ with different values to initialize and train the holistic tokenizer. Next, we train an AR generation model – Hita-L with fixed 128 holistic tokens in total. Consequently, both image reconstruction and generation quality are estimated. As listed in Table 3, when $k$=0, the token fusion module degrades to a vanilla causal transformer. In this case, the patch-level tokens can directly influence the reconstruction of their corresponding patches, creating a shortcut that bypasses the holistic tokens. As a result, the token fusion process ignores the holistic tokens, leading to its codebook collapse, as evidenced by the 2nd row in Table. 3. Once $k\textgreater 0$, the patch-level tokens can only affect the reconstruction of other patches and must also rely on the holistic tokens. In this scenario, the problem of holistic codebook collapse is avoided. Besides, the codebook utilization rate and the reconstruction quality can also be improved. Here, we also found that $k$ equals 4 achieves the optimal reconstruction and generation quality. Thus, we configure $k$ equals 4 in our experiments, by default.
@@
Pre-trained models for semantic injection. When generating holistic features, a pre-trained foundation model (e.g., DINOv2 [7]) is adopted to inject semantic information. However, the impact of using different pre-trained models to inject semantics on image reconstruction and generation quality has not yet been explored. Therefore, here we study injecting semantic-aware features from different pre-trained models, including the vision language model (e.g., CLIP [36]) or vanilla vision backbone (e.g., ResNet [14]). Here, we initialize and train a holistic tokenizer with different pre-trained foundation models with default settings. Then we train Hita-L for default 50 epochs. As shown in Table 4 , injecting DINOv2’s semantic features yields better results. Therefore, we utilize DINOv2 to perform semantic injection as default.
@@
Table 4: Comparison of different pre-trained models leveraged to perform semantic injection with advance discriminator.
What is the semantic-level of latent tokens? To analyze the semantic-level of the learned holistic tokens, we conduct linear probing according to the setting of MAE [15] on ImageNet [8] validation set. Specifically, the holistic tokens from the final causal transformer are used. Here we focus on validation instead of engaging in a comprehensive comparison with all alternative approaches.
Compared to the vanilla tokenizer VQGAN [46], we notice that Hita achieves a better accuracy, suggesting it captures semantic information. Nevertheless, it lags behind DINOv2  [7]. Because the semantic features learned by Hita are different from DINOv2’s. As shown in Fig. 2a, Hita’s primary goal is image reconstruction, focusing more on learning relatively low-level global features like texture, material, color, and other features from pixels. Nevertheless, a pre-trained model – DINOv2 integrated in Hita can inject necessary semantic-aware features to help reconstruction. Conversely, DINOv2 tends towards understanding image content, with self-supervised ViT features containing explicit information about the semantic segmentation of an image, capturing more general and high-level features.

### v4

@@ Abstract
Vanilla autoregressive image generation models generate visual tokens step-by-step, limiting their ability to capture holistic relationships among token sequences. Moreover, because most visual tokenizers map local image patches into latent tokens, global information is limited. To address this, we introduce Hita, a novel image tokenizer for autoregressive (AR) image generation. It introduces a holistic-to-local tokenization scheme with learnable holistic queries and local patch tokens. Hita incorporates two key strategies to better align with the AR generation process: 1) arranging a sequential structure with holistic tokens at the beginning, followed by patch-level tokens, and using causal attention to maintain awareness of previous tokens; and 2) adopting a lightweight fusion module before feeding the de-quantized tokens into the decoder to control information flow and prioritize holistic tokens. Extensive experiments show that Hita accelerates the training speed of AR generators and outperforms those trained with vanilla tokenizers, achieving 2.59 FID and 281.9 IS on the ImageNet benchmark. Detailed analysis of the holistic representation highlights its ability to capture global image properties, such as textures, materials, and shapes. Additionally, Hita also demonstrates effectiveness in zero-shot style transfer and image in-painting. The code is available at https://github.com/CVMI-Lab/Hita.
@@ 1 Introduction
To address these issues, we introduce Hita, a novel global-to-local image tokenizer for AR image generation. First, we propose using learnable queries to capture the holistic information from the patch-level image embeddings. Meanwhile, a pre-trained foundation model, such as DINOv2 [7], is adopted to inject semantics into the learned queries to enhance their representation. Then, we quantize the learned queries and patch-level embeddings using two separate codebooks, resulting in holistic tokens and patch-level tokens. Next, we fuse the de-quantized holistic and patch-level tokens using a novel token fusion module to reconstruct the image. This design can better prioritizes the holistic token and improving both image reconstruction and generation quality.
Once Hita is trained, a vanilla autoregressive (AR) transformer model, e.g. Llama [48], can be seamlessly incorporated to generate both holistic and image patch tokens sequentially, which can be decoded into image facilitated by Hita. During the AR generation, the AR generator first generates holistic tokens, which encapsulate global image information and serve as a prefix prompt. These holistic tokens then guide the subsequent generation of local image patch tokens, alleviating difficulties and maintaining global coherence for improved generation quality.
Our in-depth ablation experiments confirm that the learned holistic tokens effectively capture global image features, such as shape, color, texture, etc, from pixels (see Fig. 2a). This is demonstrated by effortlessly enabling style transfer by combining holistic tokens from reference images with patch-level tokens from the source image. Moreover, without task-specific training, the model can be seamlessly applied to image inpainting by encoding the given incomplete image into latent tokens, which serve as prefix sequence prompts to the AR model, requiring it to generate visual tokens for completing the image (see Fig. 2b), demonstrating the capability of holistic tokens in ensuring global consistent synthesis.
@@
In summary, our contributions are:
  • We propose a novel global-to-local image tokenizer for AR image generation that allows integration of AR generation models without modification.
  • The proposed image tokenizer emerges with new characteristics, such as zero-shot style transfer and zero-shot image in-painting.
  • Experiments on class-conditional image generation demonstrate that it accelerates convergence during training and significantly improves synthesis quality.
@@ 2 Related Work
TiTok [59] and VAR [47] are related efforts but differ from Hita in several aspects. TiTok creates compact 1D tokens with queries to reduce redundancy, while Hita uses queries to capture holistic properties for 1D holistic tokens and fuses them with 2D image patch tokens for image reconstruction. TiTok refines queries for 1D tokens tailored for reconstruction, whereas Hita extracts information from a pre-trained foundation model and patch features to capture global information. TiTok requires a complex training strategy, while Hita allows efficient single-stage training. Hita outperforms TiTok in AR generation and offers new features like zero-shot style transfer and inpainting. VAR [47] predicts global and local tokens indiscriminately, while Hita separates holistic and patch-level tokens, enriching holistic tokens with pre-trained foundation to form global representations and uses causal attention to align with the causal nature of AR models, which VAR lacks. Removing initial coarse-scale tokens seldom affects VAR’s reconstruction, indicating those tokens are not crucial.
@@ 3 Method
In this section, we first introduce the preliminaries related to image tokenizers and AR generation models. Then, we present the design of our holistic tokenizer step-by-step. Next, we discuss their characteristics.
@@ 3.1 Background
Autoregressive Image Generation. When used for visual generation, the 2D quantized map from a tokenizer is converted into a 1D sequence in a pre-defined order, e.g. raster-scan order. Then, an AR generation model autoregressively predicts the sequence of image tokens. To achieve this, AR models typically incorporate causal attention, using a prefix token sequence to predict the next, focusing on learning the dependencies between the prefix and current discrete image tokens. Considering a sequence of discrete tokens ${x_{1:N}}$, an AR model is trained on these tokens by maximizing the cross-entropy depicted in Eq. (3), where ${c}$ is class label or text embedding, ${x_{j<i}}$ serves as the prefix sequence of tokens for the current token ${x_{i}}$.
@@
During inference, the AR model initially predicts visual tokens autoregressively by sampling tokens from the learned distribution. Subsequently, the generated visual tokens are converted back into images through the decoder $\mathcal{D}(\cdot)$.
@@ 3.2 Holistic Tokenization for AR Generation
As shown in Fig. 3, Hita incorporates the following key components. First, to capture holistic properties, we introduce learnable queries that interact with image patch embeddings via attention modules. A pre-trained foundation model (e.g., DINOv2 [7]) injects semantic-level features. To align Hita’s latent space with the causal nature of AR models, we use a causal transformer to integrate holistic and patch features before quantization, embedding the sequential structure into the tokenization process for learning prefix-aware latent tokens. Details are depicted in Sec. 3.2.1. Then, the holistic and patch-level features are quantized using separate codebooks (see Sec. 3.2.2). Finally, after dequantization and during token fusion and decoding procedure, Hita first uses a fusion module with causal transformers to process these tokens in order and prioritize holistic tokens. Then they are fed into the decoder for image reconstruction (see Sec. 3.2.3).
@@ 3.2.1 Holistic Feature Extraction
We build Hita based on VQGAN [12]. Specifically, we adopt its encoder $\mathcal{E}(\cdot)$ to encode input image $I$ into local patch embeddings, and further incorporate a set of learnable queries ${Q}$ to gather holistic information from the whole input. To enhance the holistic representation, a pre-trained foundation model $\mathcal{H}(\cdot)$ like DINOv2 [7] is leveraged to inject rich semantic-level features. To achieve this, we first concatenate the learnable queries and the flattened patch-level embeddings as well as foundation model features into sequence and feed them into a simple transformer $\mathcal{E}_{\text{{trans}}}(\cdot)$ is utilized, where holistic queries ${Q}$ interact with all patch-level image embeddings and the injected semantic-level features. Since the AR model models the probability of observing the current token based on its predecessors, a latent space well-aligned with the causal nature is encouraged. Therefore, the extracted holistic and patch-level features are concatenated and reorganized in an ordered sequence, with holistic queries come first followed by the patch-level features obtained in raster-scan order. This sequence is then fed into a transformer with global features serving as prefix prior. The whole process is formulated as Eq. (4).
@@
Figure 4: Token fusion and decoding procedure
@@ 3.2.2 Separate Vector Quantization
Once both learned queries $\overline{Q}$ and patch-level images $\overline{Z}$ are produced, they are quantized with holistic quantizer $\mathcal{Q}_{H}(\cdot)$ and patch quantizer $\mathcal{Q}_{P}(\cdot)$, respectively. Then, the holistic tokens $\hat{Q}$ and patch-level $\hat{Z}$ tokens could be obtained. The the codebook significantly impacts the performance of the image tokenizer. Following [54, 46], we apply ${\ell}_{2}$-normalization to the codebook vectors, opting for a lower vector dimension and a larger codebook size to improve the reconstruction quality and codebook utilization.
@@ 3.2.3 Token Fusion and Decoding Procedure
After quantization, the holistic tokens and patch-level tokens can be fed into a standard transformer for token fusion and then fed into the decoder to reconstruct the image. However, such a design incurs codebook collapse in the holistic quantizer (3th row in Table. 3). This occurs because patch-level tokens directly impact the reconstruction of relevant image patches through the skip connection in the transformer, which makes it bypass holistic tokens (see Fig. 1a in Appendix). As a result, the token fusion module learns a trivial solution that overlooks the holistic tokens and leads to holistic codebook collapse.
@@ 3.2.4 Training objective
3.3 Autoreggresive Image Generation
Once the tokenizer is trained, a standard autoregressive (AR) transformer model, e.g. Llama [48], can be seamlessly integrated without requiring any revision to generate these latent tokens sequentially. Starting from the class/text embedding $c$, the AR model first generates holistic tokens as prefix prompts, then guides the subsequent generation of patch token sequence in the way of next-token prediction. Once all the visual tokens are generated, they undergo the token fusion and decoding procedure to be converted into pixels. Additionally, in AR models, each visual token receives a positional embedding. Since holistic tokens form a 1D sequence, we add learnable positional embeddings for them. Meanwhile, image patch tokens continue to use 2D RoPE [45] for their positional embeddings, which remains the same as LlamaGen [46]
@@ 3.4 Discussions
What do holistic tokens capture? To understand what has been learned by the holistic tokens, we perform an experiment on the learned holistic tokens. Specifically, two images are first encoded and quantized into holistic tokens and image patch tokens, respectively. Then, the holistic tokens from one image are concatenated with the patch tokens from the other. These tokens are fed through the de-quantization process and the decoder to synthesize a new image. This entire procedure is conducted using a pre-trained holistic tokenizer, without any additional model training or fine-tuning. As shown in Fig. 2a, the holistic tokens can capture global information from an image, such as texture, color, materials, shapes, and other features from the pixels. For instance, in the case of the lizard image, Hita successfully transfers the material properties of the ship from the reference image to the input image, creating a realistic mechanical lizard. More samples are shown in Appendix.
Does holistic tokenizer enhance semantic coherence? To verify whether the holistic features can enhance the global coherence of the generated content, we perform a zero-shot image in-painting evaluation. In this in-painting task, the model is forced to complete the lower half of an image given only the upper half. This setup prevents local information leakage during in-painting. We first tokenize the incompleted images into holistic and patch tokens, then feed them to the AR model, conditioned on class embeddings, to generate additional tokens. Hita-XXL is used without modifications to the model architecture or parameter tuning.
As shown in Fig. 2b, Hita effectively completes the lower part of the image based on the upper part, producing images that maintain strong semantic consistency. In contrast, outputs from LlamaGen [46] lack this consistency. For example, in the case of the fish image, Hita correctly generates a complete fish, while LlamaGen [46] mistakenly generates a “fish-bird”. These results confirm that our holistic tokenizer is semantic-aware and has a stronger ability to better generate semantically and holistically coherent examples. Furthermore, the ablation study of linear probing on holistic tokens is conducted in Sec. 4.
@@ 4.1 Setup
Hita is optimized on the ImageNet [8] training set and evaluated on the validation set. To ensure a fair comparison with LlamaGen [46], we train the tokenizer on $336\times 336$ images, which yields a comparable number of discrete tokens to it. We also adopt the tokenizer training settings from LlamaGen. Additionally, during evaluation, the test images are resized to $256\times 256$, consistent with the evaluation procedure in LlamaGen.
@@
Following the AR generation procedure in LlamaGen [46], the AR models first generate images of the same size as the image tokenizer and then resize them to $256\times 256$ for evaluation. Unless otherwise specified, all models are trained with the same settings as LlamaGen, with the entire training process lasting 50 epochs. Additionally, classifier-free guidance [18] is used in the AR generation process.
@@ 4.3 Ablation Study
First, we study the impact of learnable queries and semantic injection. Second, we analyze the selection length $k$. Next, we discuss different foundation models for semantic injection. Later, we study the semantic level of holistic tokens. We provide further ablation in the Appendix.
Learnable queries and Semantic Injection Analysis. In Hita, it incorporates a set of learnable queries and leverages semantic injection from a pre-trained model (e.g., DINOv2 [7]) to capture holistic properties. To validate their effectiveness, we conduct analysis on the learnable queries ${Q}$ and semantic injection $\mathcal{H}$. Specifically, We first train a tokenizer with only the attention modules as a baseline. Then, we gradually add learnable queries and a pre-trained foundation model (e.g., DINOv2) to this baseline. Next, we train a Hita-B AR generation model for 50 epochs to estimate the generation quality. Moreover, we conduct linear probing on the whole latent tokens after the last attention module in the baseline. As for the comparison experiments, we only perform linear probing on the holistic tokens from the last attention module. As depicted in Table. 2, simply introducing learnable queries can significantly improve image reconstruction and generation quality. The additional features from the foundation model can further enhance the reconstruction and generation quality. Besides, as indicated by the linear probing, the learnable queries can capture better semantic information than the baseline. With semantics injected, this semantic-level representation can be further enhanced. This illustrates that combining the learnable queries and semantic injection is essential for learning better holistic information, as lacking either degrades image reconstruction and generation quality.
@@
Is $k>0$ crucial? To emphasize holistic tokens and avoid its codebook collapse, Hita first builds a 1D ordered sequence with holistic tokens first and followed by the flattened patch-level tokens. Next, it is fed into a causal transformer for token fusion. Subsequently, the last $k$ holistic tokens are combined with patch-level tokens to reconstruct the image. To verify the doubt, we first initialize and train the holistic tokenizer, configuring $k$ with different values. Next, we train an AR generation model – Hita-L with fixed 128 holistic tokens in total. Consequently, both image reconstruction and generation quality are estimated. As listed in Table 3, when $k=0$, the token fusion module degrades to a vanilla causal transformer. In this case, the patch-level tokens can directly influence the reconstruction of their corresponding patches, creating a shortcut that bypasses the holistic tokens. As a result, the token fusion process ignores the holistic tokens, leading to its codebook collapse, as evidenced by the 2nd row in Table. 3. Once $k>0$, the patch-level tokens can only affect the reconstruction of other patches and must also rely on the holistic tokens. In this scenario, the problem of holistic codebook collapse is avoided. This also improves the codebook utilization and the reconstruction quality. Here, we also found that $k$ equals 4 achieves the optimal reconstruction and generation quality. Thus, we set $k$ to 4 by default in our experiments.
@@
Pre-trained models for semantic injection. When generating holistic features, a pre-trained foundation model (e.g., DINOv2 [7]) is adopted to inject semantic information. However, the impact of using different pre-trained models for semantic injection on image reconstruction and generation remains unexplored. Here we study injecting semantic-aware features from different pre-trained models, including the vision language model (e.g., CLIP [36]) or vanilla vision backbone (e.g., ResNet [14]). We initialize and train a holistic tokenizer with different pre-trained foundation models using the default settings. Then we train Hita-L for default 50 epochs. As shown in Table 4, injecting DINOv2’s semantic features yields better results. Therefore, we utilize DINOv2 to perform semantic injection as default.
@@
Table 4: Comparison of semantic injection performance with different pre-trained models.
What is the semantic-level of latent tokens? To analyze the semantic-level of the learned holistic tokens, we conduct linear probing according to the setting of MAE [15] on ImageNet [8] validation set. Specifically, the holistic tokens from the final causal transformer are used. Here we focus on estimation rather than engaging in a comprehensive comparison with all alternative approaches.
Compared to the vanilla tokenizer VQGAN [46], we notice that Hita achieves a better accuracy, suggesting it captures semantic information. Nevertheless, it lags behind DINOv2  [7]. Because the semantic features learned by Hita are different from DINOv2’s. As shown in Fig. 2a, Hita’s primary goal is image reconstruction, focusing more on learning relatively low-level global features like texture, material, color, and other features from pixels. Nevertheless, DINOv2, a pre-trained model integrated in Hita, injects necessary semantic-aware features to aid reconstruction. Conversely, DINOv2 tends towards understanding image content, with self-supervised ViT features containing explicit information about the semantic segmentation of an image, capturing more general and high-level features.